# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v3

@@ 7 Conclusion
Whilst Weyl’s foundational system does not contain quantification over exponentiation of functions as observed by [5], we have argued that it is not necessary to sacrifice exponentiation altogether to reconcile Bishop’s mathematics with Weyl’s approach if we adopt a foundation like the Minimalist Foundation.
Given that the Minimalist Foundation interprets HAω, the price to pay is to renounce the rule of unique choice as well as all the number-theoretic choice principles characteristic of Bishop’s mathematics and that real numbers à la Dedekind or Cauchy form a set. All these facts together call for a point-free development of analysis by adopting the topological methods advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45].
Hence, provided that a point-free reformulation of classical analysis is viable, as hinted in [7, 42, 43, 21], MF  promises to be a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity.
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterpart, as established for MF in [33], or at least whether they remain predicative after the addition of the law of excluded middle. A positive answer to either of these questions would provide a more expressive constructive foundation capable of formalizing general results of point-free mathematics as those presented in [12, 37, 15, 47], by keeping the compatibility both with Bishop’s constructivism and with Weyl’s classical predicativity, as in the case of MF.

### v4

@@ 7 Conclusion
Whilst Weyl’s foundational system does not contain quantification over exponentiation of functions as observed by [DBLP:journals/bsl/Avron20], we have argued that it is not necessary to sacrifice exponentiation altogether to reconcile Bishop’s mathematics with Weyl’s approach if we adopt a foundation like the Minimalist Foundation.
Given that the Minimalist Foundation interprets HAω, the price to pay is to renounce the rule of unique choice as well as all the number-theoretic choice principles characteristic of Bishop’s mathematics and that real numbers à la Dedekind or Cauchy form a set. All these facts together call for a point-free development of analysis by adopting the topological methods advocated by Martin-Löf and Sambin with the introduction of Formal Topology [sambin1987].
Hence, provided that a point-free reformulation of classical analysis is viable, as hinted in [CN95, Palmgren05, pal07, kawai2023], MF promises to be a natural crossroads between Bishop’s constructivism and Weyl’s classical predicativity.
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extensions of MF with inductive and coinductive definitions in [MMR21, MMR22, MS2023] are still equiconsistent with their classical counterpart, as established for MF in [MSdoubleneg], or at least whether they remain predicative after the addition of the law of excluded middle. A positive answer to either of these questions would provide a more expressive constructive foundation capable of formalizing general results of point-free mathematics as those presented in [CSSV03, mv04, curicind, sambinbook], by keeping the compatibility both with Bishop’s constructivism and with Weyl’s classical predicativity, as in the case of MF.